OpenStudy (anonymous):

A water trough is 15 m long and has a cross-sectional shape of an equilateral triangle with a side length of 2 m. It is being filled at a rate of 0.6 m3/min. How fast is the water level rising when the water is 0.2 m deep?

OpenStudy (amistre64):

regardless of side length; you need to find away to equate a change in height with a change in volume i believe

OpenStudy (amistre64):

V = basearea*15; gotta determine the area for an equilat i spose dV/dt = .6

OpenStudy (amistre64):

base = s height = s sin(60) V = 15 s^2 sqrt(3)/4 dV = 30sqrt(3)/4 *s *ds/dt perhaps?

OpenStudy (amistre64):

find dh/dt height = s sin(60) s = h/sin(60) = 2h/sqrt(3) ds/dt = 2/sqrt(3) dh/dt dh/dt = sqrt(3)/2 * ds/dt

OpenStudy (amistre64):

\[\frac{dV}{dt} = \frac{30s\sqrt{3}}{4}\frac{ds}{dt}\] \[.6 = \frac{30s\sqrt{3}}{4}\frac{2}{\sqrt{3}}\frac{dh}{dt}\] \[.6 = \frac{30s}{2}\frac{dh}{dt}\] \[\frac{.6(2)}{30s} =\frac{dh}{dt}\] \[\frac{.12}{30s} =\frac{dh}{dt}\] when we know s at time (t) then we know dh/dt

OpenStudy (amistre64):

well, we dont need to know t :) but the problem allows us to determine s when the water is at any given height

OpenStudy (amistre64):

1-sqrt(3)-2 = b/2-h-s h = .2 h = sqrt(3) scalar is .2/sqrt(3) 2(.2) ----- = s sqrt(3) \[\frac{dh}{dt}=\frac{.12\sqrt{3}}{2(.2)(30)}\] \[\frac{dh}{dt}=\frac{.12\sqrt{3}}{.4(30)}\] \[\frac{dh}{dt}=\frac{12\sqrt{3}}{40(30)}\] \[\frac{dh}{dt}=\frac{\sqrt{3}}{10(10)}\] \[\frac{dh}{dt}=\frac{\sqrt{3}}{100}\] maybe?

OpenStudy (amistre64):

by the chain rule: \[\frac{dh}{dt}=\frac{dh}{ds}\frac{ds}{dV}\frac{dV}{dt}\] dV = .6 ................................................... V = 15 s^2 sqrt(3)/4 1 = 30s sqrt(3)/4 ds ds = 4/(30s sqrt(3)) ; s = .4/sqrt(3) = 4/30(.4) = 1/30(.1) = 1/3 ................................................... h = s sin(60) dh = sin(60) = sqrt(3)/2 \[\frac{dh}{dt}=\frac{\sqrt{3}*1*6}{2*3*10}\] \[\frac{dh}{dt}=\frac{\sqrt{3}}{10}\] ........................................... which may be a better result